can you help me solve by completing the square?
x^2-6x-3=0

Answer by stanbon(75887)

solve by completing the square?
x^2-6x-3=0
Rearrange:
x^2-6x+? = 3+?
Now, complete the square
x^2-6x+(6/2)^2 = 3+(6/2)^2
Factor the left side and evaluate the right side:
(x-3)^2 = 12
Take the square root of both sides to get:
x-3 = 2sqrt3 or x-3=-2sqrt3
Solve for x to get:
x=3+2sqrt3 or x=3-2sqrt3
============
Cheers,
Stan H.